# A Neuro-Symbolic AI Approach to Personal Health Risk Assessment and Immune Age Characterisation using Common Blood Markers

### v7

@@ 2.5 NHANES Age Prediction
The sex groups were separated by health status as determined by binning the immune scores and ages by the healthy range thresholds indicated in Table  1, which are as follows: Red Blood Cell (RBC) count (between 3.8 to 5.8), and the following healthy ranges for the multivariate measure: normal Mean Corpuscular Haemoglobin (MCH) (27-33), Haematocrit (HCT) (37-47), lymphocytes (LYMP) (1.0-4.0), and neutrophils (NEUT) (1.8-7.5). Above or below these healthy count ranges, individuals were binned into the unhealthy groups. Further, the error rates analysis for the age prediction between the actual ages and immune-score computed ages was performed using these binned subsets.
@@
The zero counts were removed from the NHANES data prior to all analysis. However, for the quantitative analysis, sex was replaced by 1 for males and 2 to denote females. Table 2 highlights sex-specific differences in immune analyte values, for reference.
@@ 3 Results
We have introduced a set of scores for patient triaging purposes tested against a large dataset based on numerical distance and a colour scheme. The most relevant part of the score is its single numerical value ranging from 0 to 10, capturing the distance of all other test values from population or personal reference values. Results show that such triage is statistically significant even in the face of noise, as in the NHANES database (containing 100K unfiltered surveys and tests), where disease is self-reported by interviewees who may have had the disease at the time of the survey or at some prior point in time. Filtering the score by test results also makes for a significant improvement, showing that it will perform better under less noisy conditions.
Figure 3: Immune score space: The $x$ axis represents the deviation of all the analytes from the expected value, while the $y$ axis measures the total number of analytes. A negative standard deviation indicates that the value is lower than expected and a positive one indicates a higher than expected value. The cut-off seen on the negative $x$ axis shows the fact that real-life values for analytes have a strict lower bound but, technically, have no upper bounds. For instance, it is not possible to have a negative white blood cell count.
The behaviour of the numerical value is consistent through varying numbers of analytes and scales linearly with respect to deviation from the norm, and this graph explores the behaviour of the immune score over synthetically generated patient values as a function of number of analytes and divergence from healthy reference values.
Figure 4: The behaviour of NCCIS over synthetically generated patients with differing values and a deviation from the norm with regard to the respective expected values. The $x$ axis measures the stated analyte, while the $y$ axis measures the total number of analytes. In contrast to the behaviour of the immune score (Fig. 3), NCCISS is a non-linear function that displays a step-wise behaviour by design, while remaining independent of the number of analytes.
@@
Table 4: Differences in NIS mean values. All comparisons were performed using a one-way ANOVA.
Figure 5: Distribution of the Normalised Immune Score (NIS) from 0 to 10 ($y$ axis) among adult individuals in the NHANES 2000-2014 database. The Healthy cohort is defined as those individuals with all analytes within the normal ranges as defined by the NHS (1); others comprise the Unhealthy set. The Self rep. healthy cohort comprises individuals who self-reported health of 3 (”Good”) or less. The number of individuals in each cohort is shown in table 4.
@@
Table 5: Differences in NCCIS mean values. All comparisons were performed using a one-way ANOVA. Effect of sample size was analysed in Table 6 for statistical significance and practical applicability.
@@
Table 6: Cohen’s d effect sizes of combined blood parameters in the NIS immune Score, and Immune Age comparing healthy and unhealthy groups. Effect sizes highlight significant differences between the two groups. P-values were computed using an independent Welch’s t-test, which compares the means of two groups. Results are statistically significant, suggesting robust differences. The Immune Score (SCORE) indicates a discriminatory distinction between the two groups for practical purposes with SCORE, the immune score, and immune age having a statistical power above 1.00 (alpha 0.05) suggesting a low risk of false discovery. The Immune Age by Health status shows a strong Cohen’s d (2.28), indicating it is a highly robust measure for differentiating health status and predicting blood parameter-based health risks with precision. In contrast, the weaker effect size by Sex (0.11) suggests that sex is unlikely to be a confounding factor, reinforcing the accuracy of the Immune Age prediction based on health status.
Figure 6: Distribution of the Normalised Colour-Coded Immune Score (NCCIS) from 0 to 10 ($y$ axis) among adult individuals in the NHANES 2000-2018 database. Individuals are defined as Healthy if all their analytes are within the normal ranges as defined by the NHS (1) and Unhealthy otherwise. Self rep. healthy comprises individuals who self-reported health of 3 (“Good”) or less. The number of individuals in each cohort is shown in table 5. One explanation for the bimodality of some of these plots is the nature of the source data as self-reported, where respondents may or may not have been ill at the time of giving their answers, or from the application of the separating cutoff value from NIS. In either case, both score-separating capabilities are highly statistically significant. The separation of groups of self-reported data suggests that they could be separated from those that were actually not healthy at the time of testing. In contrast, those self-reported not healthy were indeed clinically not healthy according to their results and the score.
@@
In Table 7, we provide examples of CCIS and its correspondent NCCIS values for a sample of real-life and artificial cases. The second column of the table shows the NIS, for purposes of comparison. It is worth noting that these examples were generated using the NHS values for normal healthy subjects, and based on the provisional assumption that the distance between the upper and normal ranges is twice the standard deviation. The examples were taken from [9] as they were intended as very preliminary tests. We are aware that their source clearly states they are meant only for teaching purposes. Here they are used for purposes of illustration only, and the development of the NIS and NCCIS is not dependent on them. In Table 6, we provide Cohen’s d values suggesting practical applicability of the score, making it clinically relevant for triaging purposes even at the level of individual classification and prediction.
@@
One can observe how the NCCIS values meet the requirements we set in advance of their definition. We tested the NIS and NCCIS against real data from the National Health and Nutrition Examination Survey 2003-2016 (NHANES), provided in [32]. One measure of success was the ability to show how accurately the score discriminates between healthy subjects and patients suffering from various diagnosed illnesses.
@@
To this end, we examined thousands of cases from [32] and calculated the distribution, mean values, and standard deviations for each of the analytes (see Fig. 7. We used the means and standard deviations from the data to define the metrics for both the NIS and NCCIS. Finally, we calculated the NIS and NCCIS for thousands of cases in the NHANES database to find individual scores. The resulting values were grouped as belonging to healthy individuals or to a selected list of common diseases. Furthermore, Figure 8 demonstrates how machine learning predictions can dynamically adjust the expected range of analyte values, detecting stable trends while identifying outliers, thus enhancing precision in monitoring patient health.
@@
In order to maximise the discriminatory power of both scores, we tried different alternative combinations of means, normal ranges and standard deviations, either taking these values from NHS values as of early 2022 or inferring them from the NHANES database (Fig. 7). At the end of the day, we settled for (1) NHS normal healthy reference values and means and standard deviations for the NIS; (2) means and standard deviations calculated from the NHANES database, and normal ranges from the NHS for the NCCIS. NHANES means and standard deviations are shown in Table 3.
@@
Figure 5 illustrates the distribution of NIS values across various (self-declared) conditions for individuals in the NHANES database. The table 4 presents the number of data points used, along with the respective ANOVA results to test statistical significance. In principle, smaller sample sizes in one group could lead to unstable variability in p-values. However, the consistent statistical significance across all comparisons and the large effect sizes shown in Table 6 indicate that the results are robust. Furthermore, Cohen’s d is based on the entire dataset, using group means, variances, and sample sizes to compute the pooled standard deviation, thereby accounting for sample sizes in the reported effect sizes. The large effect size of the immune age indicates it significantly distinguishes healthy from unhealthy individuals, reflecting its statistical and practical significance in health status discrimination and prediction. It must be noted that the self-reported conditions were not independently confirmed. Moreover, the survey participants did not distinguish between current or past diagnoses. This could be expected to introduce some noise, as some currently healthy people will be labelled as having a condition and some people with conditions will not have been diagnosed. We hope to improve accuracy by filtering the data or by adding other data sources in the future.
@@
Nevertheless, the NIS was able to discriminate between healthy and non-healthy individuals, as most healthy individuals are clustered around very low NIS values. In contrast, different conditions produced higher NIS values on average.
@@
The method is shown to be valid for predicting immune profiles/trends (variation over time) observed in diverse patients when considering socio-clinical metadata parameters such as age and sex, thereby suggesting their clinical applicability for purposes such as triaging and fast screening and extensions to deviant clinical cases in prospective studies. In Fig. 6, it can be seen that reported medical conditions produced a higher average NCCIS value, but also induced some anomalous clustering. Furthermore, the distribution of scores was even wider than that of the NIS. We thus concluded that the NCCIS did enhance the insights already gleaned from the NIS and the CCIS.
@@
Table 7: Different tested scores perform differently according to each score’s definition.
@@
Statistical tests assessed the relationship among the three individualised scores, NIS, NCCIS, and CCIS, across various health conditions. Table 8 presents significant statistical associations between these scores. The Kruskal-Wallis test, a non-parametric alternative to ANOVA was performed due to the ordinal nature of the CCIS score. The significant p-value (0.0053) indicates that there are statistically significant differences in the distribution of NIS, NCCIS, and CCIS scores across the conditions. A Pearson correlation assessed the linear relationship between NIS and NCCIS. The strong positive correlation coefficient (0.89) with a highly significant p-value (¡ 0.0001) determined from a two-tailed independent samples t-test suggests a robust linear association between these two scores. Lastly, a Wilcoxon Signed-Rank test compared paired differences between NIS and NCCIS. The significant result (p = 0.018) indicates that there are significant differences between these paired scores.
3.1 Linearisation and Separability
The scores introduced here are not intended to be used as a diagnostic tool on their own and can only quantify abnormality in terms of deviation from healthy reference (absolute or adaptable) values. The score is sensitive to out-of-range values and increases its value or changes its colour as a function of how removed values are from lower and upper bounds of reference values according to the number of standard deviations from the medians, but it cannot quantify diseases or conditions. The score indicates how far the bulk of all markers are from normal (healthy) reference values, the median, and an interval determined by published reference values for specific demographic or health conditions. Medians can be derived from empirical data (Fig. 1).
3.2 Adaptive Reference Values
Figure 7: Personalised reference values: Statistical inference without smart outlier detection. In contrast to the smart outlier detection technology, simple statistical regression towards the observed values results in skewed inferred ranges.
Figure 8: Smart outlier detection example. The machine learning technology can adapt the range of expected values for all analytes, showing stable trends while also detecting abnormal values or trends. In the image, we can show the result of machine learning inference of the ranges of normal values for a patient while also detecting abnormal values.
4 Immune Age Estimation
The immune system plays an important role in protecting against infections and in maintaining human health. Blood is the medium in which most immune cells circulate through the body.
Ageing is associated with complex changes and dysregulation of cellular processes. Nine tentative hallmarks that represent common denominators of ageing in different organisms have been described before[28].
It has been postulated that inflammation plays a critical role in regulating physiological ageing. Inflammatory components of the immune system are often chronically elevated in aged individuals, a phenomenon that has been termed “inflammageing” [15, 17, 16, 18, 22, 4].
However, the dynamics of this process at the individual level have not been characterised, hindering quantification of an individual’s ‘immune age’.
[4] used multiple ‘omics’ technologies to capture population- and individual-level changes in the human immune system in individuals of different ages sampled longitudinally over a nine-year period. They observed high inter-individual variability in the rates of change of cellular frequencies that was dictated by the individual’s baseline values, allowing identification of steady-state levels toward which a cell subset converged and the ordered convergence of multiple cell subsets toward older adult homeostasis.
In [21], an IMM-AGE score was described that captured an individual’s immune-ageing process. The IMM-AGE score correlated with age, yet it captured additional metrics such as cell-cytokine response better than chronological age.
Here we present a method where a single blood test, as in a complete full blood count of 13 blood analytes, affords insight into the immune age as compared to chronological age. Chronological age simply indicates the age of the individual, as determined by the length of time someone has lived or has been living since their birth. This term should not be confused in the context of other literature, such as age estimates predicted by DNA methylation patterns from the Horvath algorithm.
4.1 Immune Score Over time
@@
We hypothesised that as the human body aged the immune score would vary over time. In particular, we expected to see an upwards trajectory in direct relation to age.
@@
To test the hypothesis, we selected from the dataset two cohorts comprising all the data entries of individuals with ages ranging between 20 and 90 years, separating them by sex (male or female). These cohorts were used to build two immune spaces by computing the respective mean and standard deviations for each of the thirteen analytes. Afterward, for each sex, we selected seven cohorts of adults aged 20 to 25, 25 to 35, 35 to 45, 45 to 55, 55 to 65, 65 to 75, and 75 to 85 years of age, making for a total of 14 cohorts. Finally, we computed the immune score for each entry. The mean immune age for each of the cohorts is shown in Fig. 9.
@@
Figure 9: (a) The mean immune age as a function of age for NHANES data. The reference ranges were computed over the population aged 20 to 85 for males and females, respectively. Each of the data points represents the mean immune score for the cohort [age5,age+5][age-5,age+5]; for example, 50 represents the immune score for the population aged 45 to 55 years. (b)The distribution of immune age as a function of age. While a correlation trend is visible (especially for males), across sex the correlation is weak because healthy and non-healthy cohorts are confounded at the source given that both and the patient’s actual current unhealthy states are in the same category.
@@
The results show significantly different behaviour over time for male and female cohorts. For subjects over the age of 30, it is evident that the male population presents an upward trajectory with respect to age for the mean immune score, as well as an increase in the 50th percentile and variance. The immune score decreases slightly for males between 20 and 25 year old. However, the trend in females is slowed down (compared to males) between ages 40 and 50 and reversed between ages 50 and 60. It remains to be investigated whether some of the analytes involved are confounded with contradictory trends from other physiological differences, such as menopause, which occurs exactly between the ages where the slow-down and reversal in the age score trend for females is detected. It is known and has been reported that pregnancy and menopause produce a higher and more progressive increase in red cell counts, haematocrit levels and increased mean cell volume (MCV) as well as haemoglobin concentrations [25, 6, 11, 31, 2].
4.2 Discounting Confounding Factors in Immune Age

### v9

@@ 2.2 NHANES Age Prediction
The sex groups were separated by health status determined by dividing immune scores and ages by the healthy range thresholds indicated in Table  S1 (see Sup. Inf.), which are as follows: Red Blood Cell (RBC) count (between 3.8 to 5.8), and the following healthy ranges for the multivariate measure: normal Mean Corpuscular Haemoglobin (MCH) (27-33), Haematocrit (HCT) (37-47), lymphocytes (LYMP) (1.0-4.0), and neutrophils (NEUT) (1.8-7.5). Above or below these healthy count ranges, the individuals were categorised into unhealthy groups. In addition, error rates analysis for age prediction between actual ages and immune index computed ages was performed using these binned subsets.
@@
The zero counts were removed from the NHANES data prior to all analysis. However, for the quantitative analysis, sex was replaced by 1 for males and 2 to denote females. Table S2 (see Sup. Inf.) highlights sex-specific differences in immune analyte values, for reference.
2.3 Immune Age Estimation
The immune system plays an important role in protecting against infections and maintaining human health. Blood is the medium in which most immune cells circulate through the body.
Ageing is associated with complex changes and dysregulation of cellular processes. Nine tentative hallmarks that represent common denominators of ageing in different organisms have been described [41].
It has been postulated that inflammation plays a critical role in the regulation of physiological ageing. Inflammatory components of the immune system are often chronically elevated in aged individuals, a phenomenon that has been termed “inflammageing” [33, 34, 35, 36, 37, 17].
However, the dynamics of this process at the individual level have not been characterised, hindering quantification of an individual’s ‘immune age’. In Ref. [17], various ‘omics’ technologies are used to capture population- and individual-level changes in the human immune system in individuals of different ages sampled longitudinally over a nine-year period. They observed high inter-individual variability in the rates of change of cellular frequencies that was dictated by the individual’s baseline values, allowing identification of steady-state levels toward which a cell subset converged and the ordered convergence of multiple cell subsets toward older adult homeostasis. In [38], an IMM-AGE index was described that captured an individual’s immune-ageing process. The IMM-AGE index was correlated with age, yet it captured additional metrics such as cell-cytokine response better than chronological age.
Here, we present a method where a single blood test, as in a complete full blood count of 13 blood analytes, provides insight into immune age compared to chronological age. Chronological age simply indicates the age of the individual, determined by the length of time someone has lived or has lived since their birth. This term should not be confused with the context of other literature, such as age estimates predicted by DNA methylation patterns from the Horvath algorithm.
2.4 Machine Learning Methodology
NHANES data on blood analytes (CBC markers) and NIS immune scores between patients were separated by health status for a binary classification problem: healthy vs. unhealthy. The data were pre-processed by imputing the missing NaN values using the mean values for each feature. The data set was then divided into training and test sets for classification, model building, and evaluation. Test sizes of 0.2 and 0.5 were used with a 10-fold cross-validation for analysis of performance metrics. Five statistical machine learning (ML) classifiers were used for initial assessment: Random Forest (RF), Support Vector Machine (SVM), K-Nearest Neighbours (KNN), KMeans clustering, and Gaussian Mixture Model. The latter two clustering methods delivered poorer performances in the data features, thus were removed from the screening. Each model was trained in the training set and evaluated in the discarded test set. Feature importance was analysed for the RF, SVM, and KNN models to identify the most predictive features. Pairwise analysis was also performed to inspect feature relationships. A correlation analysis was performed to identify any co-linear features before saving the results.
Two artificial neural network (ANN) architectures were explored as more flexible artificial intelligence (AI) tools. ANN models tested in binary health classification by analytes and immune scores included a feedforward neural network (FFNN) (also referred to as a multilayer perceptron MLP). The MLP consists of an input layer, one or more hidden layers, and an output layer. The model uses ReLU (Rectified Linear Unit) activation functions for the hidden layers and a sigmoid activation function for the output layer, which is typical for binary classification tasks.
Additionally, in order to make the ANN model more robust, we incorporated a more complex layered architecture with a Transformer network. The Keras and TensorFlow Python packages were used for the development of the neural network model. Statistical metrics such as ROC curves (area under the curve- AUC), sensitivity, specificity, mean square error (MSE) and classification accuracy were computed to evaluate the predictive performance of the ML algorithms. A description of the algorithms used is given in the Sup. Inf.
2.5 Further Validation of the Indexes with the UK Biobank database
The UK Biobank is a long-term prospective nation-wide study that houses the deidentified biological samples and health-related data of half a million people. Between 2006 and 2010 volunteers aged 40 to 69 years were recruited from all over Great Britain and consented to share their health data and to be followed for at least 30 years thereafter; with the objective of allowing scientific discoveries to be made in the prevention, diagnosis, and treatment of disease [20, 22]. We tested the NIS index (which is the NCCIS index without the warning colouring component, see Supp. Inf.), to separate large disease categories in the UK Biobank. The database does not contain a clean healthy control group, meaning that all participants had potentially some reported health problem, making our separation results more relevant given that we had to create a Rest group different from the large disease categories identified as a noisy control group.
@@ 3 Results
The results show that routine and common markers from CBC tests can be exploited for personalised healthcare and precision medicine. That CBC markers can be grouped into numerical scores that can be used for triaging purposes even from noisy self-reported sources, such as in the NHANES database (containing 100 000 unfiltered surveys and tests). Filtering by normal and abnormal results improves the detection of health-related patterns and significantly improves the classification of health status, demonstrating that the approach performs better under less noisy conditions, as expected.
In contrast to the behaviour of the immune index, NCCISS is a non-linear function that displays a step-wise behaviour by design S2 (Sup Inf.), while remaining independent of the number of analytes.
To assess group differences in immune scores, we first conducted Shapiro–Wilk tests to evaluate normality. Due to violations of normality assumptions and the ordinal nature of the NCCIS score, we employed the non-parametric Kruskal–Wallis test to assess overall group differences. All comparisons were statistically significant (p << 0.0001). To evaluate differences between specific group pairs, we used Mann–Whitney U tests with Bonferroni correction for multiple comparisons, retaining significance across all contrasts.
@@
Table 1: Differences in NIS mean values. Statistical comparisons were performed using the Kruskal–Wallis test, followed by Mann–Whitney U tests with Bonferroni correction for pairwise group comparisons.
@@
Table 2: Differences in NCCIS mean values. Kruskal–Wallis tests revealed significant group differences (p << 0.0001), followed by Bonferroni-corrected Mann–Whitney U tests for pairwise comparisons. Effect sizes are discussed in Table 3.
@@
Table 3: Cohen’s d effect sizes of combined blood parameters in the NIS immune index and Immune Age, comparing healthy and unhealthy groups. While Cohen’s d is robust to normality violations, we additionally report p-values using Welch’s t-test. Results demonstrate strong practical and statistical significance of the immune index.
The Immune Age by Health Status shows a strong Cohen’s d (2.28) 3, indicating that it is a strong and robust measure to differentiate health status and predict health risks based on blood parameters with precision. In contrast, the weaker effect size by sex (0.11) suggests that sex is unlikely to be a confounding factor, reinforcing the precision of the Immune Age prediction based on health status.
Figure 1: Distribution of the Normalised Colour-Coded Immune index (NCCIS) from 0 to 10 ($y$ axis) among adult individuals in the NHANES 2000-2018 database. Individuals are defined as Healthy if all their analytes are within the normal ranges as defined by the NHS (S1 (see Sup. Inf.) and Unhealthy otherwise. Self rep. healthy comprises individuals who self-reported health of 3 (“Good”) or less. HIV is an example of chronic infection while all the diseases can fall in the category of immune and blood-related conditions.
@@
Table 4 shows examples of CCIS and its corresponding NCCIS values for a sample of real-life and artificial cases. The second column of the table shows the NIS, for purposes of comparison. It is worth noting that these examples were generated using the NHS values for normal healthy subjects and based on the provisional assumption that the distance between the upper and normal ranges is twice the standard deviation. The examples were taken from [21] as they were intended as very preliminary tests. We are aware that their source clearly states that they are meant only for teaching purposes. Here, they are used for purposes of illustration only, and the development of the NIS and NCCIS is not dependent on them. In Table 3, we provide Cohen’s d values that suggest practical applicability of the index, making it clinically relevant for triaging purposes even at the level of individual classification and prediction.
@@
One can observe how the NCCIS values meet the requirements we have set in advance of their definition. We tested NIS and NCCIS against real data from the 2003-2016 National Health and Nutrition Examination Survey (NHANES), provided in [16]. One measure of success was the ability to show how accurately the index discriminates between healthy subjects and patients suffering from various diagnosed diseases.
@@
We calculated the distribution, mean values, and standard deviations for each of the analytes (see Fig. S2 (top)). We used means and standard deviations from the data to define the metrics for both the NIS and NCCIS. Finally, we calculate NIS and NCCIS for thousands of cases in the NHANES database to find individual scores. The resulting values were grouped as belonging to healthy individuals or to a selected list of common diseases. Furthermore, Fig. S2(Bottom) (see Suppl. Info) shows how machine learning predictions can dynamically adjust the expected range of analyte values, detecting stable trends while identifying outliers, thus improving precision in monitoring patient health.
@@
To maximise the discriminatory power of both scores, we tried different alternative combinations of means, normal ranges, and standard deviations, taking either these values from NHS values as of early 2022 or inferring them from the NHANES database  S2(top). We settled for (1) NHS normal healthy reference values and means and standard deviations for the NIS; (2) means and standard deviations calculated from the NHANES database, and normal ranges of NHS for NCCIS. NHANES means and standard deviations are shown in Table S3 (see Sup. Inf.)
@@
The table 1 presents the number of data points used, together with the respective Mann Whitney U test results to test pairwise statistical significance. In principle, smaller sample sizes in one group could lead to unstable variability in p-values. However, the consistent statistical significance of all comparisons and the large effect sizes shown in Table 3 indicate that the results are robust. Furthermore, Cohen’s d is based on the entire dataset, using group means, variances, and sample sizes to compute the pooled standard deviation, thereby accounting for sample sizes in the reported effect sizes. The large effect size of immune age indicates that it significantly distinguishes healthy from unhealthy individuals, reflecting its statistical and practical significance in discrimination and prediction of health status. It should be noted that the self-reported conditions were not independently confirmed. In addition, the survey participants did not distinguish between current or past diagnoses. This could be expected to introduce some noise, as some currently healthy people will be labelled as having a condition and some people with conditions will not have been diagnosed. We hope to improve accuracy by filtering the data or by adding other data sources in the future.
@@
However, the NIS was able to discriminate between healthy and nonhealthy individuals, since most healthy individuals are clustered around very low NIS values. In contrast, different conditions produced higher NIS values on average.
@@
The method is shown to be valid for predicting immune profiles/trends (variation over time) observed in diverse patient groups when considering socio-clinical metadata parameters such as age and sex, thereby suggesting their clinical applicability for purposes such as triaging and fast screening and extensions to deviant clinical cases in prospective studies. In Fig. 1, it is shown that reported medical conditions produced a higher average NCCIS value, but also induced some anomalous clustering. Furthermore, the distribution of scores was even wider than that of the NIS. We therefore conclude that the NCCIS did enhance the insights already gleaned from the NIS and the CCIS. Separation of groups of self-reported data suggests that they could be separated from those that were not actually healthy at the time of testing. In contrast, those who self-reported not being healthy were indeed clinically not healthy according to their results and the index. The number of individuals in each cohort is shown in table 2. One explanation for the bimodality of some of these plots is the nature of the source data as self-reported, where respondents may or may not have been ill at the time of giving their answers, or from the application of the separating cutoff value from NIS.
@@
Table 4: Different tested scores perform differently according to each index’s definition.
@@
Statistical tests assessed the relationship among the three individualised scores, NIS, NCCIS, and CCIS, for various health conditions. Table 5 presents significant statistical associations between these scores. The Kruskal-Wallis test, a non-parametric alternative to ANOVA was performed due to the ordinal nature of the CCIS index. The significant p-value (<<0.001) indicates that there are statistically significant differences in the distribution of the NIS, NCCIS, and CCIS scores in the conditions. A Pearson correlation assessed the linear relationship between NIS and NCCIS. The strong positive correlation coefficient (0.89) with a highly significant p-value (<< 0.0001) determined from a two-tailed independent samples t-test suggests a robust linear association between these two scores. Lastly, a Wilcoxon Signed-Rank test compared paired differences between NIS and NCCIS. The significant result (p = 0.018) indicates that there are significant differences between these paired scores and therefore each can capture different aspects of blood data.
3.1 Adaptive Index Behaviour Over Time
In 2, we tested the behaviour of the NIS index over time by simulating patients evolving in different directions with outliers. We show that the index can be used to learn setpoints or baselines of patients over time based on the behaviour of the index flagging for changes even within the normal values of the population.
Figure 2: Simulation of a process of learning personalised reference values based on NIS index behavious. The index is used to adapt the range of expected values for all analytes over time, showing stable trends while also detecting abnormal values from a personal previously established setpoint over time. Dashed lines are personalised adaptive reference values for two types of adaptation with and without outlier detection.
@@ 3.1 Adaptive Index Behaviour Over Time
We also hypothesised that as the patient aged, the immune index would vary over time. In particular, we expected to see an upward trajectory in direct relation to age.
@@
To test the hypothesis, we selected from the dataset two cohorts comprising all the data entries of individuals with ages ranging between 20 and 90 years, separating them by sex (male or female). These cohorts were used to build two immune spaces by computing the respective mean and standard deviations for each of the 13 analytes. Afterward, for each sex, we selected seven cohorts of adults aged 20 to 25, 25 to 35, 35 to 45, 45 to 55, 55 to 65, 65 to 75, and 75 to 85 years of age, making a total of 14 cohorts. Finally, we computed the immune index for each entry. The mean immune age for each of the cohorts is shown in Fig. 3.
@@
Figure 3: (a) The mean immune age as a function of age for NHANES data. The reference ranges were computed over the population aged 20 to 85 for males and females, respectively. Each of the data points represents the mean immune index for the cohort [age5,age+5][age-5,age+5]; for example, 50 represents the immune index for the population aged 45 to 55 years. (b)The distribution of immune age as a function of age. While a correlation trend is visible (especially for males), across sex the correlation is weak because healthy and non-healthy cohorts are confounded at the source given that both and the patient’s actual current unhealthy states are in the same category.
@@
The results show significantly different behaviour over time for male and female cohorts. For subjects over the age of 30, it is evident that the male population presents an upward trajectory with respect to age for the mean immune index, as well as an increase in the 50th percentile and variance. The immune index decreases slightly for males between 20 and 25 year old. However, the trend in women slows down (compared to men) between ages 40 and 50 and reverses between ages 50 and 60. It remains to investigate whether some of the analytes involved are confounded with contradictory trends from other physiological differences, such as menopause, which occurs exactly between the ages when the slowdown and reversal of the age index trend for women is detected. Pregnancy and menopause are known to produce a more progressive and higher increase in red blood cell counts, hemocrit levels, and increased mean cell volume (MCV), as well as hemoglobin concentrations [40, 18, 23, 43, 12].
3.2 Discounting Confounding Factors from Age